A girl has a number of ducks costing $10 each and three times as many chickens costing $6 each. If the total cost of ducks and chickens is $420, find the number of chickens the girl has.
Answer by haileytucki(390)

A girl has a number of ducks costing $10 each and three times as many chickens costing $6 each. If the total cost of ducks and chickens is $420, find the number of chickens the girl has.
x(10)+3x(6)=420
Multiply x by each term inside the parentheses.
10x+3x(6)=420
Multiply 3x by each term inside the parentheses.
10x+18x=420
Since 10x and 18x are like terms, add 18x to 10x to get 28x.
28x=420
Divide each term in the equation by 28.
(28x)/(28)=(420)/(28)
Simplify the left-hand side of the equation by canceling the common factors.
x=(420)/(28)
Simplify the right-hand side of the equation by simplifying each term.
x=15
Now, if x=15 then 15*10=150 ducks